OpenStudy (whereisthefloor):

Jeff wants to rent a boat. He has to pay a fixed base cost plus a daily rate for renting the boat. The table shows the amount of money, y, in dollars, that Jeff has to pay for renting the boat for x days:

OpenStudy (mathstudent55):

Let us think this through. The cost of the boat rental is y. y = a daily cost + a fixed cost

OpenStudy (mathstudent55):

Notice what happens on day zero. Even if he rents the boat for no days, it is already $3. That means $3 does not depend on the number of days. $3 is the fixed cost. We now have y = daily cost + 3

OpenStudy (mathstudent55):

Now let's see what happens on each day. If he rents for 1 day, he pays $8. We know that $3 is the fixed amount, so we subtract $8 - $3 = $5. The daily charge for 1 day is$5. Let's look at the second day just to confirm the daily charge. If he rents for 2 days, he pays $13. We subtract the fixed charge to find the daily charge for 2 days: $13 - $3 = $10 $10/2 = $5 Once again, we see he pays $5 per day.

OpenStudy (mathstudent55):

Yes, B and C both have a fixed charge of $8 which is not correct. Once again, we have y = daily charge + 3 What is the daily charge in terms of x, the number of days? This is just the daily charge: If he rents for 1 day, he pays 5 * 1 = 5 If he rents for 2 days, he pays 5 * 2 = 10 If he rents for 3 days, he pays 5 * 3 = 15 If he rents for x days, he pays 5 * x = 5x The daily charge is 5x. What is y, then? (daily charge + fixed charge)

OpenStudy (mathstudent55):

No. I'm not asking for a number. I didn't tell you how many days he's renting. I'm asking for the correct expression for the cost. The rental has a fixed cost of $3 and a daily cost of 5x. y = daily cost + fixed cost What is y?

OpenStudy (mathstudent55):